what is -3 + 2b =-9
Add 3 to both sides:
3 + -3 +2b = -9 + 3
That cancels the 3's on the one side and you get 2b by itself.
2b = -9 + 3
2b = -6
Divide both sides by 2, so you can have the b by itself
2b/2 = -6/2
b = -3


Answer by elima(1433)

what is -3 + 2b =-9
Add 3 to both sides:
3 + -3 +2b = -9 + 3
That cancels the 3's on the one side and you get 2b by itself.
2b = -9 + 3
2b = -6
Divide both sides by 2, so you can have the b by itself
2b/2 = -6/2
b = -3
Check:
-3 + 2(-3) = -9
-3 + -6 = -9